CDP

CDP 0.771 (06.09.95) CDP - A CDPlayer for SCSI-2 CD-ROM-Drives with a MUI-Interface.

Copyright c_{◯} 1994 Andreas "andreash" Harrenberg

1 Copyright

1.1 Programmstatus

CDP 0.771 (06.09.95) Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994 Andreas "andreash" Harrenberg

This program is freely distributable, but copyrighted by me. This means you may freely copy it as long as you do not charge more than a nominal fee for copying. This program may be put on PD disks, especially on Fred Fish’s AmigaLibDisks and Fred’s CDROMs. If you want to distribute this program you MUST keep this document with it. The program and the documents must be distributed in their original unmodified form.

This program cannot be used for commercial purposes without written permission by the author. The author can not be made responsible for any damage which is caused by using this program.

Send comments, suggestions or bug reports, and the gifts to:

Andreas Harrenberg, Hubertusstr. 55 D-52064 Aachen, Germany Phone: +49 (0)241 - 23784

INTERNET: andreash@pool.informatik.rwth-aachen.de INTERNET: andreash@mowgli.fido.de INTERNET: Andreas.Harrenberg@post.rwth-aachen.de FIDO: 2:2452/107.7@fidonet (Light) FIDO: 2:242/7.7@fidonet (Classic)

2 What is CDP?

2.1 Programdescription

CDP is a simple CD Player for SCSI-II CD-Rom drives. It needs MUI (at least V2.3) for its GUI-Interface.

In the GUI you are able to Show/Hide things like Skip-Buttons or Number-Buttons. There is also an optional Window containing the songs of the CD. (Songsfile aren’t yet implemented, sorry!)

The main playerwindow can be very small, if you only display the CD-Display and the Start- /Stop-Button.

CDP is in a very first stage of development, so don’t expect too much from this release. The main purpose of this release is to get feedback from users.

3 What is needed?

3.1 Requirements on Hard- and Software

- Kickstart >= 2.0

- Workbench >= 2.0

- MUI >= 2.3

- SCSI-II CD-ROM

- SCSI-Interface capable of sending scsi-direct-commands.

4 Installation

4.1 How to install CDP?

The installation of CDP is really simple. As there are yet no Songfiles implemented it is sufficient to simply copy the files CDP and CDP.info to an directory of your choice.

If you i.E. want to copy the programm to Work:CDROM-Soft, and the Archiv was extracted to RAM: type the following in the shell:

copy ram:CDP/bin/CDP ram:CDP/bin/CDP.info work:CDROM-Soft/

From Workbench it is simply a drag and drop. Drag the Icon with the programm over a dircetory of your choice and ready!

There is also a simple installerskript (Install English) which copy the programm and the manual.

5 CLI-Arguments

5.1 Options from Shell

TEMPLATE:

-D=DEVICE/K,-U=UNIT/K,-S=SONGDIR/K,-DF=DISPLAYFONT/K,-BF=BIGDISPLAYFONT/K
-D=DEVICE/K

Voreingestellt ist <scsi.device> Devicename of the scsi-interface. Note that this name is case sensitiv! The suffix <.device> is optional, CDP will add <.device> for you. Default is <scsi.device>

-U=UNIT/K

The unit-number of your CD-Rom drive. Default is 5.

-S=SONGDIR/K

Directory for the songfiles. (SONGFILES NOT YET IMPLEMENTED)

-DF=DISPLAYFONT/K

Small (normal) font for the display, defaults to topaz.font/8.

-BF=BIGDISPLAYGFONT/K

Big font for the display, defaults to topaz.font/16.

6 WB-Tooltypes

6.1 Options from Workbench

The (long) Section 5.1 [CLI-Arguments], page 5 are recognized as tooltypes.

DEVICE= UNIT= SONGDIR= DISPLAYFONT= BIGDISPLAYFONT=

7 GUI

7.1 The Userinterface

After the first start CDP opens two windows, the first is the control-window with the display and the second displays all songs in a listview.

The listview contains information about start- and playtime of each track, the type of the track (data/audio) and the title. The title is set to Nr. x, but as soon as the songfiles are implemented you can edit and set this to the real songnames. A doubleclick or a mui-select (cursor & return) will start playing the selected song. You can open/close this window (MENU: GUI/SONGS...), and the state of the window will be saved, so after the next start of CDP the window may not be open.

The control-window diplays the cd-title (if known by the programm) in the first textobject, the display itself shows track/index/playtime. In future version there will be more (configurable) infos such as totalplaytime etc.

The next group are numbergadgets from 1 to 9. This buttons can be used to start the songs from 1 to 9. Sorry, two digit numbers are not yet implemented. You can show/hide this group from the menu (GUI/NUMBERS...), the state will be saved after quiting CDP.

The skip-section consist of search/index and skip, search jumps 10 seconds forward/back, index jumps to next/previous index, and skip plays the next/previous song. This group can also made invisible (MENU: GUI/SKIPS...), the state is saved.

The index-skip only worked, if the CD-Rom drive is capable of this feature, if not either nothing will happen, or the song is played from the start again. The CD300 (Apple) doesn’t has a correct index-function, this will be emulated in a later version. As a result of this, the index-buttons will be disabled if a CD300 is detected.

Ok, last but not least, play/pause/stop. These gadgets are alwalys visible, as long as you don’t iconfy the whole application.

8 Future

8.1 What to do

- Songfiles...

- new display

- Index for ALL drives ?

- programms

9 Thanks

9.1 Thanks are going to…

- kmel, Klaus Melchior for his help with MUI - Stefan Stuntz for his MUI

- Eric Boëz for his nice letter

- and all I forgot to mention ...